# Need Help? Designing Proactive AI Assistants for Programming

### v1

@@ 5.4 Experimental Platform
The study was conducted online and asynchronously at the participant’s own time. The interface we used to deploy the study is an extension of the open-source platform RealHumanEval introduced in [18]. Conducting the study on the web allows for easy assignment of participants to experimental conditions, the duration of the study, and the tasks participants solve. Finally, the web interface reduces installation issues or any potential incompatibility. While the web interface is not the traditional IDE that a user may typically use for their day-to-day development, it is reminiscent of online coding platforms (e.g., Leetcode) that participants are generally familiar with. The editor in our web interface is also the same Monaco editor as in Visual Studio Code, a popular IDE. We also added a check in the post-study form to make sure that participants did not have any issues using the interface.
5.5 Procedure
@@
Before participating in the study, each participant filled out a consent form. The study has been approved by our institution’s review board (IRB). We adopt a between-subject setup to compare the three proactive chat variants and a within-subjects setup to compare each proactive variant to the non-proactive baseline. This means each participant will interact with both a proactive assistant and the baseline chat condition. We randomize the order in which participants interact with either the proactive assistant or the baseline chat assistant. The total time the participants spent coding for the study was 40 minutes, with participants spending 20 minutes in each condition.
5.5.1 Onboarding
@@
Before participants can access the proactive assistant, they are provided with a short tutorial that describes how to interact with the proactive suggestions. The tutorial highlights how the proactive assistant will provide a high-level summary of suggestions that can be expanded upon and how to ask follow-up questions. The tutorial also warns participants that the suggested implementations may not always be correct and to verify the suggestions carefully. In the Appendix, we provide a copy of the onboarding instructions that participants had access to.
5.5.2 Coding Tasks
@@
In our study, participants engage in 20-minute sessions focusing on two types of programming problems designed to reflect real-world engineering scenarios. The first type of problem aims to test system-building skills. For instance, participants are tasked with enhancing an “online store” class object implementation. Given an initial starter code, they must implement additional features specified through natural language instructions, brainstorm a new feature, debug potential issues, and write tests for these new features. The second type of problem challenges participants to work with unfamiliar packages, simulating situations where engineers must quickly adapt to new tools. An example of this is manipulating a dataframe using various NumPy functions. Unlike previous studies that primarily assessed puzzle-solving abilities through Leetcode-style problems [14, 18], our approach aims to evaluate skills more directly relevant to practical software engineering tasks. To avoid confounding task difficulty or type with the helpfulness of proactive assistants, participants will start with the same task type in both proactive and baseline conditions, though the tasks themselves are randomized. In the Appendix, we provide descriptions for each task in the study.
5.5.3 Post-task survey
@@ 5.6 Measurements
  1. 1.
    Telemetry. Our interface logs all user behavior from coding to using the chat assistant: anytime the code is updated by the user, the interface saves the updated code. Further, any chat messages and associated responses are logged. Finally, any interactions with the proactive assistant, including expanding suggestions, accepting a suggestion, and requesting suggestions are all recorded. Using the collected telemetry data, we compute the the number of sub-tasks completed. We can also obtain fine-grained interaction behavior to identify common interaction patterns with proactive suggestions.
  2. 2.
@@ 5.7 Analysis Approach
We measure the effect of proactive assistants on user productivity (RQ1) by computing the number of sub-tasks completed across each of the four coding tasks. We build a linear model which incorporates the experimental condition, which is either baseline or one of the proactive models, and the coding problem as fixed effects. We also qualitatively inspect whether the code written contains automated test cases, which is supplementary measure of well-written code. To understand the effect of proactive assistants on user experience (RQ2), we measure participant ratings of the baseline and proactive assistant. We then computed a binary measure of whether participants preferred the proactive assistant to the baseline and ran an ANOVA with Tukey HSD. We also analyze participant’s written justifications to interpret the quantitative results. For all tests, the threshold for statistical significance was $\alpha=0.05$. To begin to understand how participants use the different proactive assistants (RQ3), we measure how often participants used the various components of the proactive assistants as logged in the telemetry (i.e., expanding, previewing, accepting, copying, and requesting suggestions). We consider this portion of the results an exploratory analysis to identify trends that distinguish the proactive conditions and understand the effect of different design decisions on user interactions. We present the results of our study in the following section.
6 Results
6.1 RQ1: Proactive Assistants Generally Improve User Productivity
@@
Figure 3: Percentage of sub-tasks completed correctly. Comparing baseline chat to proactive assistants across the four tasks, where Task 1 and Task 2 are system-building questions and Task 3 and 4 are ones where participants work with new packages and functionality. While performance varied by task, we observed that all variants of proactive assistants tended to increase the number of test cases passed compared to the baseline chat assistant across the board.
Number of sub-tasks completed.
@@
When comparing the number of sub-tasks completed by participants, we find that on average participants with proactive assistants are more productive with a baseline chat assistant: we observe a $12.1\%\pm 5.1\%$, $18\%\pm 5.8\%$, and $11.6\%\pm 5.0\%$ increase in the percentage of test cases passed for Suggest and Preview, Suggest, and Persistent Suggest respectively compared to baseline. Figure 3 provides a more granular view by task. We find that the improvements in the number of test cases over the baseline chat assistant are significant across all proactive variants, where $p=0.01$ for Suggest and Preview, $p=0.002$ for Suggest, and $p=0.02$ for Persistent Suggest.
Number of test cases written.
@@
In addition to the increase in the number of test cases passed, we also observe an increase in the amount and quality of test cases written. Note that writing tests was part of the instructions provided to users. Focusing on Task Type B, where it is important to test how different components of the designed system work together, we annotate whether the user code contains test cases. We find that only 13.3% do in the baseline setting while 33.3% do in the proactive settings. We believe that providing proactive suggestions can help create a virtuous cycle where the suggestions generate tests for code the participant has written to surface issues in their code, thus increasing the number of test cases passed.
6.2 RQ2: User Experience with Proactive Assistants Varies by Implementation.
@@
Figure 4: Comparing participant perception of proactive versus baseline. User experience is important to the adoption of proactive assistants. We compare how often participants preferred a proactive assistant to the baseline variant in each condition. We find participants generally preferred the proactive variant only in the Suggest and Suggest and Preview conditions.
User preferences across proactive conditions.
@@
While we observe that proactive assistants generally improve user productivity, user experience and perceptions of the proactive assistant vary across the conditions. As shown in Figure 4, we measure whether participants preferred the proactive assistant over the baseline non-proactive assistant. We find that the vast majority of participants prefer the proactive variant over the baseline for both the Suggest and Suggest and Preview proactive assistants (90% and 80.1% of participants respectively). In contrast, less than half of participants (45%) in the Persistent Suggest proactive condition preferred having the proactive assistant. We find that both Suggest and Suggest and Preview variants are statistically different than the Persistent Suggest variant ($p=0.005$ and $p=0.03$ respectively). On average, participants had a neutral view of the baseline assistant in terms of usefulness. As such, participants viewed both the Suggest and Suggest and Preview variants on average as net beneficial.
@@
Participants in the Persistent Suggest condition tended to not prefer the proactive assistant because they often found the assistant to be “distracting” and “annoying”. One participant noted that “the non-proactive chat assistant was best because it didn’t interrupt what I was doing.” and another participant “found it annoying because it distracted [them] from working with the non-proactive answers.” In contrast, participants in the Suggest and Suggest and Preview conditions tended to have a positive view of the proactive assistant. One participant noted that the reason they preferred the proactive assistant was because “non-proactive was pretty useless because it didn’t have any of the context regarding what I was trying to do or the task; I could only really use it for pure syntax like what I would usually search StackOverflow for. Proactive was better because it had more context.” and another stated that “I really wasn’t sure what to ask for with the non-proactive chat” since the baseline chat “required manual input to generate advice”. Next, we analyze how user perception translates into the actual usage of the proactive assistant.
6.3 RQ3: How Do Participants Use Different Proactive Assistants?
@@
To better understand the effect of different design decisions on how participants use proactive assistants, we break down user interactions with proactive suggestions by condition. First, we compare Persistent Suggest and Suggest conditions to explore the effect of suggestion timing. Second, we compare Suggest and Suggest and Preview conditions to explore the effect of the level of proactivity.
@@
Figure 5: Comparing interactions with different proactive conditions. Visualizing sampled participant trajectories from coding with different proactive assistants. We denote “expanding” a suggestion with E, “accepting” a suggestion with A, “previewing” a suggestion with P, and “copying” a suggestion with C. In the Persistent Suggestion condition, participants tend to expand suggestions without utilizing them further. In contrast, participants tend to copy expanded suggestions in the Suggest condition while participants in the Suggest and Preview condition tend to preview suggestions.
6.3.1 The effect of suggestion timing.
@@
6.3.2 The effect of level of proactive assistant involvement.
The ability to preview and incorporate suggestions into code changed how users interacted with the respective proactive assistants. For example, after expanding a suggestion, further interactions with suggestions were more than twice as likely to be through the preview functionality rather than copying code and integrating it into the editor by themselves ($3.1\pm 1.5$ in the Suggest and Preview condition compared to $0.8\pm 0.5$ Suggest condition)—an example of such an interaction pattern is shown in Figure 5. This may also have led to an increase in manual requests for suggestions, which happened more often in the Suggest and Preview condition, where participants made on average $2.9\pm 0.9$ manual requests in the Suggest and Preview compared to an average $1.7\pm 0.7$ requests in the Suggest condition. These results suggest that participants may gravitate towards options that reduce their manual efforts to make changes, which include utilizing the preview functionality.
7 Discussion
7.1 Design Implications For Proactive Coding Assistants
7.1.1 Revisiting design considerations
@@
(1) Support efficient evaluation. We observe that participants generally felt that the proactive assistant provided suggestions that had “clear answers while giving specific examples” and allowed them to quickly “get a baseline understanding of the problem.”, which provides evidence that the suggestions were presented in a way that allowed participants to understand and use them efficiently. Further, we believe the suggestions were not too distracting; another participant noted that the suggestions helped to “pinpoint specific issues while not being so obnoxious [they] couldn’t ignore it to focus on other tasks.”
@@
(4) Time services based on context. Aside from the Persistent Suggest condition where participants felt suggestions appeared too often (as discussed in Section 6.2), participants had positive perceptions of the proactive suggestion timings, noting that the proactive assistant was “always ready” and comparing the proactive assistant to “a second set of eyes” or “a second person solving [the problem] for me by my side.” Additionally, multiple participants suggested that the timing of suggestions can even enable them to “catch issues early” or “fix errors early on so that they don’t require refactoring the entire codebase later”. Our results across proactive conditions demonstrate that appropriately timed suggestions can be very helpful to users, however inappropriately timed suggestions may be perceived as more of a nuisance.
(5) Incorporating user feedback. Interestingly, this was the only design consideration that received mixed feedback from participants. Some participants believed “the ability to accept or reject the code it suggested was really nice” and “intuitive” to use while others felt like they were “unnecessary” and “interacting with them distracts me from the task.” Future work may consider incorporating user feedback implicitly rather than requiring users to accept or delete suggestions.
7.1.2 Expanding design considerations

### v2

@@ 5.4. Experimental Platform
The study was conducted online and asynchronously at the participant’s own time. The interface we used to deploy the study is an extension of the open-source platform RealHumanEval introduced in (Mozannar et al., 2024c). Conducting the study on the web allows for easy assignment of participants to experimental conditions, the duration of the study, and the tasks participants solve. Finally, the web interface reduces installation issues or any potential incompatibility. While the web interface is not the traditional IDE that a user may typically use for their day-to-day development, it is reminiscent of online coding platforms (e.g., Leetcode) that participants are generally familiar with. The editor in our web interface is also the same Monaco editor as in Visual Studio Code, a popular IDE. We also added a check in the post-study form to make sure that participants did not have any issues using the interface.
5.5. Procedure
@@
Before participating in the study, each participant filled out a consent form. The study has been approved by our institution’s review board (IRB). We adopt a between-subject setup to compare the three proactive chat variants and a within-subjects setup to compare each proactive variant to the non-proactive baseline. This means each participant will interact with both a proactive assistant and the baseline chat condition. We randomize the order in which participants interact with either a proactive assistant or the baseline chat assistant. The total amount of time the participant spent coding for the study was 40 minutes, with participants spending 20 minutes in each condition.
5.5.1. Onboarding
@@
Before participants can access the proactive assistant, they are provided with a short tutorial that describes how to interact with the proactive suggestions. The tutorial highlights how the proactive assistant will provide a high-level summary of suggestions that can be expanded upon and how to ask follow-up questions. The tutorial also warns participants that the suggested implementations may not always be correct and to carefully verify the suggestions. In the Appendix, we provide a copy of the onboarding instructions that participants had access to.
5.5.2. Coding Tasks
@@
In our study, participants engage in 20-minute sessions focusing on two types of programming problems designed to reflect real-world engineering scenarios. The first type of problem aims to test system-building skills. For instance, participants are tasked with enhancing an “online store” class object implementation. Given an initial starter code, they must implement additional features specified through natural language instructions, brainstorm a new feature, debug potential issues, and write tests for these new features. The second type of problem challenges participants to work with unfamiliar packages, simulating situations where engineers must quickly adapt to new tools. An example of this is manipulating a dataframe using various NumPy functions. Unlike previous studies that primarily assessed puzzle-solving abilities through Leetcode-style problems (Vaithilingam et al., 2022; Mozannar et al., 2024c), our approach aims to evaluate skills more directly relevant to practical software engineering tasks. In the Appendix, we provide task descriptions for each of the tasks used in the study.
The problem types are randomized across conditions and in both proactive/baseline conditions to avoid confounding task difficulty or type with the helpfulness of proactive assistants. Since each participant experienced both the baseline and proactive conditions, we distributed the 4 tasks in a symmetrical manner. This means that if the participant started with a system-building task, they would see the other system-building task when the proactive assistant was switched on or off. This scheme also ensures a relatively similar number of task types in both proactive and baseline conditions. Participants decide when they are satisfied with their current attempt on the given task and would like to move on to another question. As such, we had a second question of the other task type queued up for the participant to work on for the remainder of the time within that condition. In sum, each participant attempted two tasks of the same type across the baseline and proactive conditions, and additionally one or two of the remaining tasks of the different task type when time permitted.
5.5.3. Post-task survey
@@ 5.6. Measurements
  1. (1)
    Telemetry. Our interface logs all user behavior from coding to using the chat assistant: anytime the code is updated by the user, the interface saves the updated code. Further, any chat messages and associated responses are logged. Finally, any interactions with the proactive assistant, including expanding suggestions, accepting a suggestion, and requesting suggestions are all recorded. Using the collected telemetry data, we compute the the number of sub-tasks completed. We can also obtain fine-grained interaction behavior to identify common interaction patterns.
  2. (2)
@@ 5.7. Analysis Approach
To measure the effect of proactive assistants on user productivity (RQ1), we computed the number of sub-tasks completed. We build a linear model which incorporates the experimental condition, which is either baseline or one of the proactive models, and the coding problem as fixed effects. We also qualitatively inspect whether the code written contains automated test cases. To understand the effect of proactive assistants on user experience (RQ2), we measure participant ratings of the baseline and proactive assistant. We then computed a binary measure of whether participants preferred the proactive assistant to the baseline and ran an ANOVA with Tukey HSD. We also analyze participant justifications to interpret the quantitative results. For all tests, the threshold for statistical significance was $\alpha=0.05$. To begin to understand how participants use the different proactive assistants (RQ3), we measure the frequency of interactions with the various components of the proactive assistants as logged in the telemetry. We consider this portion of the results an exploratory analysis to identify trends that distinguish proactive conditions. We present the results of our study in the following section.
6. Results
6.1. RQ1: Proactive Assistants Generally Improve User Productivity
@@
Figure 4. Percentage of sub-tasks completed correctly. Comparing baseline chat to proactive assistants across the four tasks, where Task 1 and Task 2 are system-building questions and Task 3 and 4 are ones where participants work with new packages and functionality. We report average performance and standard error. While performance varied by task, we observed that all variants of proactive assistants tended to increase the number of test cases passed compared to the baseline chat assistant across the board.
[…]
Percentage of test cases passed. Comparing baseline chat to proactive assistants across the four tasks, where Task 1 and Task 2 are system-building questions and Task 3 and 4 are ones where participants work with new packages and functionality. We report average performance and standard error. While performance varied by task, we observed that all variants of proactive assistants tended to increase the number of test cases passed compared to the baseline chat assistant across the board.
@@
Number of sub-tasks completed. When comparing the number of sub-tasks completed by participants, we find that on average participants with proactive assistants are more productive with a baseline chat assistant: we observe a $12.1\%\pm 5.1\%$, $18\%\pm 5.8\%$, and $11.6\%\pm 5.0\%$ increase in the percentage of test cases passed for Suggest and Preview, Suggest, and Persistent Suggest respectively compared to baseline. Figure 4 provides a more granular view by task. We find that the improvements in the number of test cases are significant across all proactive variants, where $p=0.01$ for Suggest and Preview, $p=0.002$ for Suggest: $p=0.002$, and $p=0.02$ for Persistent Suggest. We explore whether different groups of participants benefit from proactive suggestions differently (Appendix C): we do not observe any significant differences based on Python expertise or AI tool usage frequency but do observe a significant difference based on gender (i.e., women tend to benefit more from proactive assistance).
@@
Number of test cases written. In addition to the increase in the number of test cases passed, we also observe an increase in the amount and quality of test cases written. Note that writing tests was part of the instructions provided to users. Focusing on Task Type B, where it is important to test how different components of the designed system work together, we annotate user code for whether they incorporated test cases and find that only 13.3% do in the baseline setting while 33.3% do in the proactive settings. We believe that providing proactive suggestions can help create a virtuous cycle where including more test cases helps participants surface issues in their code, thus also increasing the number of test cases passed.
6.2. RQ2: User Experience with Proactive Assistants Varies by Implementation.
@@
Figure 5. Comparing participant perception of proactive versus baseline. User experience is important to the adoption of proactive assistants. We compare how often participants preferred a proactive assistant to the baseline variant in each condition and find that participants generally preferred the proactive variant only in the Suggest and Suggest and Preview conditions.
[…]
Comparing participant perception of proactive versus baseline. User experience is important to the adoption of proactive assistants. We compare how often participants preferred a proactive assistant to the baseline variant in each condition and find that participants generally preferred the proactive variant only in the Suggest and Suggest and Preview conditions.
@@
User preferences across proactive conditions. While we observe generally uniform benefits of proactive assistants across the different conditions in terms of user productivity, we see more variation in terms of user experience across the conditions. As shown in Figure 5, we measure whether participants preferred the proactive assistant over the baseline non-proactive assistant. We find that the vast majority of participants prefer the proactive variant over the baseline for both the Suggest and Suggest and Preview proactive assistants (90% and 80% of participants respectively). In contrast, less than half of participants (47%) in the Persistent Suggest proactive condition preferred having the proactive assistant. We find that both Suggest and Suggest and Preview variants are statistically different than the Persistent Suggest variant ($p=0.005$ and $p=0.03$ respectively). On average, participants had a neutral view of the baseline assistant in terms of usefulness. As such, participants viewed both the Suggest and Suggest and Preview variants on average as net beneficial.
@@
Participants in the Persistent Suggest condition tended to not prefer the proactive assistant because they often found the assistant to be “distracting” and “annoying”. One participant noted that “the non-proactive chat assistant was best because it didn’t interrupt what I was doing.” and another participant “found it annoying because it distracted [them] from working with the non-proactive answers.” In contrast, participants in the Suggest and Suggest and Preview conditions tended to have a positive view of the suggestions provided by the proactive assistant compared to the baseline. One participant noted that they preferred the “non-proactive was pretty useless because it didn’t have any of the context regarding what I was trying to do or the task; I could only really use it for pure syntax like what I would usually search StackOverflow for. Proactive was better because it had more context.” and another stated that “I really wasn’t sure what to ask for with the non-proactive chat” since the baseline chat “required manual input to generate advice”.
Further, we identified two reasons for why participants did not prefer proactive suggestions across proactive conditions. The first reason was concerns about irrelevant suggestions, where participants thought “it was hard to clearly understand what their use cases were” and “they did not seem useful to the particular tasks”. Interestingly, participants were more concerned about relevancy, and did not mention the correctness at all, potentially due to the limited scope of tasks considered in the study. The second reason was that participants were generally unfamiliar with a proactive assistant compared to the baseline chat, saying “I was already familiar with how to code with chatbot assistants.”. Further, they already knew how to use the baseline chat well as “it feels similar to gpt-4o and copilot” and they knew what scope of tasks the chat was helpful for (e.g., “very small and simple cases”). Next, we analyze how user perception translates into the actual usage of the proactive assistant.
6.3. RQ3: How Do Participants Use Different Proactive Assistants?
@@
To better understand the effect of different design decisions on how participants use proactive assistants, we break down user interactions with proactive suggestions by condition. First, we compare Persistent Suggest and Suggest conditions to explore the effect of suggestion timing. Second, we compare Suggest and Suggest and Preview conditions to explore the effect of the level of proactivity. Finally, we investigate what kinds of suggestions participants tend to accept or reject. The following discussion should be viewed as an exploratory analysis; a later addition of statistical tests via Student t-tests did not show significant differences between conditions.
@@
Figure 6. Comparing interactions with different proactive conditions. Visualizing sampled participant trajectories from coding with different proactive assistants. We denote “expanding” a suggestion with E, “accepting” a suggestion with A, and “copying” a suggestion with C—all of which are available in every condition, and “previewing” a suggestion with P, which is only available in the Suggest and Preview condition. In the Persistent Suggestion condition, participants tend to expand suggestions without utilizing them further. In contrast, participants tend to copy expanded suggestions in the Suggest condition while participants in the Suggest and Preview condition tend to preview suggestions.
[…]
Comparing interactions with different proactive conditions. Visualizing sampled participant trajectories from coding with different proactive assistants. We denote “expanding” a suggestion with E, “accepting” a suggestion with A, and “copying” a suggestion with C—all of which are available in every condition, and “previewing” a suggestion with P, which is only available in the Suggest and Preview condition. In the Persistent Suggestion condition, participants tend to expand suggestions without utilizing them further. In contrast, participants tend to copy expanded suggestions in the Suggest condition while participants in the Suggest and Preview condition tend to preview suggestions.
6.3.1. The effect of suggestion timing.
@@
6.3.2. The effect of level of proactive assistant involvement.
The ability to preview and incorporate suggestions into code led to a change in user interactions with proactive assistants. For example, after expanding a suggestion, further interactions with suggestions were more than twice as likely to be through the preview functionality rather than copying code and integrating it into the editor by themselves ($3.1\pm 1.5$ in the Suggest and Preview condition compared to $0.8\pm 0.5$ Suggest condition)—an example of such an interaction pattern is shown in Figure 6. This may also have led to an increase in manual requests for suggestions, which happened more often in the Suggest and Preview condition, where participants made on average $2.9\pm 0.9$ manual requests in the Suggest and Preview compared to an average $1.7\pm 0.7$ requests in the Suggest condition. These results suggest that participants may gravitate towards options that reduce their manual efforts to make changes, which include utilizing the preview functionality.
6.3.3. What kinds of suggestions do participants tend to accept?
While most proactive interactions fall under expanding, copying, or previewing suggestions, as discussed in Section 6.3.1 and 6.3.2, there were a handful of interactions where participants accepted or rejected suggestions. We identified a total of 75 interactions where 69 were accepts and 6 rejects. The most commonly types of accepted suggestions were those on brainstorming new functionality and debugging (latent errors), with 18 occurrences each, while the least likely accepted suggestions were explaining existing code and pointers to documentation, with only 1 occurrence each. These trends suggest that participants are likely to accept more “actionable” suggestions, rather than ones that may be merely informative. Interestingly, improving efficiency and modularity suggestions were most commonly rejected suggestion types (with 5 out of 6 occurrences and the other being a completing unfinished code suggestion). This behavior may have been influenced by the particular study design which emphasized code correctness rather than optimized or well-written code, suggesting that different suggestions may be preferred by programmer’s with different goals. Full results are provided in Appendix C.
7. Discussion
7.1. Design Implications For Proactive Coding Assistants
7.1.1. Revisiting design considerations
@@
(1) Support efficient evaluation. We observe that participants generally felt that the proactive assistant provided suggestions that had “clear answers while giving specific examples” and allowed them to quickly “get a baseline understanding of the problem.”, which provides evidence that the suggestions were presented in a way that allowed participants to efficiently understand them and determine whether to use them. Further, the suggestions were not too distracting; another participant noted that the suggestions helped to “pinpoint specific issues while not being so obnoxious [they] couldn’t ignore it to focus on other tasks.”
@@
(4) Incorporating user feedback. Interestingly, this was the only design consideration that received mixed feedback from participants. Some participants believed “the ability to accept or reject the code it suggested was really nice” and “intuitive” to use while others felt like they were “unnecessary” and “interacting with them distracts me from the task.” Future work may consider incorporating user feedback implicitly rather than requiring users to accept or delete suggestions. Related to our findings from Section 6.3.3, different kinds of suggestions may be preferred at different points of a programmer’s workflow.
(5) Time services based on context. Aside from the Persistent Suggest condition where participants felt suggestions popped up too often (as discussed in Section 6.2), participants had positive perceptions of the proactive suggestion timings, noting that the proactive assistant was “always ready” and comparing the proactive assistant to “a second set of eyes” or “a second person solving [the problem] for me by my side.” Additionally, multiple participants suggested that the timing of suggestions can even enable them to “catch issues early” or “fix errors early on so that they don’t require refactoring the entire codebase later”.
7.1.2. Expanding design considerations